# Optimal Procurement Design: A Reduced Form Approach

### v1

@@ 2.2 Transforming the problem: A reduced form approach
Say that the collection of interim allocation probabilities $\{P_{s}\}_{s=1}^{n}$, where $P_{s}:[0,1]\to[0,1]$ for each $s$, is implementable if there exists a collection of allocation probabilities $\{p_{s}\}_{s=1}^{n}$ satisfying (F) that induces $\{P_{s}\}_{s=1}^{n}$ as its interim allocations; that is, for each $s=1,\ldots,n$ and all $q_{s}\in[0,1]$,
@@
Since the sellers are symmetric, it is without loss to restrict attention to symmetric interim allocations; thus, we can drop the subscript $s$ from $P_{s}$ and $q_{s}$, and write $P$ and $q$ instead. Consequently, the objective further reduces to
@@
Consider the quantile $s=F(q)$, we define
$\widetilde{P}(s):=P(F^{-1}(s))$
as the quantile interim allocation. Let $\widetilde{P}^{*}(s):=(1-s)^{n-1}$; it is not difficult to see that $\widetilde{P}^{*}(\cdot)$ is the quantile interim allocation of the allocation rule that always procures from the seller with the lowest quality.
Border’s (1991) celebrated theorem characterizes the set of implementable interim allocations. Lemma 2 translates Border’s condition into majorization terminology. [^4] To the best of our knowledge, Gershkov, Goeree, Kushnir, Moldovanu, and Shi (2013) is the first paper that connects Border’s condition to majorization (see their Footnote 4). We omit the proof of Lemma 2 since it can be proved by slightly modifying the proof of, for example, Theorem 1 in Hart and Reny (2015) or Theorem 3 in Kleiner et al. (2021).
@@
A decreasing interim allocation rule $P$ is implementable if and only if the associated quantile interim allocation $\widetilde{P}(s)$ is weakly majorized by $\widetilde{P}^{*}$. [^5] We use “increasing” and “decreasing” in the weak sense: “strict” will be added whenever needed.
@@
$h_{\gamma}(q):=v(q)-q-\gamma\frac{F(q)}{f(q)}$
@@
$g(q):=h_{1}(q)=v(q)-q-\frac{F(q)}{f(q)}$
denote the buyer’s virtual surplus. By Lemma 2, the quantile interim allocation of a direct mechanism that maximizes the weighted average can be found by solving
@@
$\Omega_{w}(\widetilde{P}^{*}):=\left\{\widetilde{P}\in L^{1}(0,1):\widetilde{P}~{}\text{ is decreasing and }~{}\widetilde{P}\prec_{w}\widetilde{P}^{*}\right\}.$
A solution to problem (6) exists: $\Omega_{w}(\widetilde{P}^{*})$ is compact by the Helly’s selection theorem, [^6] By “compact” we mean compact in the $L^{1}$ norm topology. For a complete proof of this fact, see the proof of Proposition 1 in Kleiner et al. (2021). and therefore the constraint set
$\Omega_{w}(\widetilde{P}^{*})\cap\left\{\widetilde{P}\in L^{1}(0,1):\int_{0}^{1}g(F^{-1}(s))\widetilde{P}(s)\,\mathrm{d}s\geq 0\right\}$
is the intersection of a closed set and a compact set and hence also compact. The argument is completed by noting that the objective function is continuous.
In particular, the buyer’s optimal interim allocation should solve the following problem:
@@
3 Buyer’s optimal procurement mechanisms
For cleaner intuition and simpler notation, in this section we study a special case of our problem: the buyer’s expected payoff maximization problem, where the welfare weight on the buyer is $\gamma=1$. In Section 3.1, we solve the problem (7), and in Section 3.2 we identify a trading mechanism that implements the solution we find. All proofs in this section are relegated to Appendix B.
3.1 Buyer’s optimal interim allocation
To ensure that the monotonicity constraint in problem (7) holds, the ironing technique (Myerson, 1981; Toikka, 2011) may be required. Define $\widetilde{g}(s):=g(F^{-1}(s))$ as the quantile virtual surplus, and let
$G(s):=\int_{0}^{s}\widetilde{g}(x)\mathrm{d}x.$
Let $\overline{G}$ be the concave hull of $G$: $\overline{G}(x):=\sup\{y:(x,y)\in\mathrm{co}(G)\}$, where $\mathrm{co}(G)$ is the convex hull of the graph of $G$. Equivalently, $\overline{G}$ is the pointwise smallest upper semicontinuous and concave function that lies above $G$ (see, for example, Kamenica and Gentzkow, 2011). Call $\overline{g}:=\overline{G}^{\prime}$ the ironed quantile virtual surplus. [^7] Because $\overline{G}$ is concave, it is differentiable almost everywhere. At points where it is not differentiable, we define $\overline{g}$ as the right derivative. Adopting the convention that $\sup\varnothing=0$, define
$\overline{S}:=\sup\{s\in[0,1]:\overline{g}(s)\geq 0\};$ (8)
our assumptions on $v$ and $F$ guarantee that $\overline{S}$ is well-defined.
@@ 3.1 Buyer’s optimal interim allocation
Let $\{\left[\underline{s}_{i},\bar{s}_{i}\right)\}_{i\in\mathcal{I}}$ denote a collection of disjoint intervals with $\left[\underline{s}_{i},\bar{s}_{i}\right)\subseteq[0,\overline{S}]$ for each $i\in\mathcal{I}$, such that
  • $\overline{G}$ is affine on $\left[\underline{s}_{i},\bar{s}_{i}\right)$ for each $i\in\mathcal{I}$, and
  • $\overline{G}=G$ on $[0,\overline{S}]\big{/}\bigcup_{i\in\mathcal{I}}\left[\underline{s}_{i},\overline{s}_{i}\right)$.
@@
$\hat{P}(s)=\begin{cases}(1-s)^{n-1}&\text{ if }s\in[0,\overline{S}]\big{/}\bigcup_{i\in\mathcal{I}}\left[\underline{s}_{i},\overline{s}_{i}\right),\\ \frac{\int_{\underline{s}_{i}}^{\overline{s}_{i}}(1-s)^{n-1}\,\mathrm{d}s}{\overline{s}_{i}-\underline{s}_{i}}&\text{ if }s\in\left[\underline{s}_{i},\overline{s}_{i}\right),\\ 0&\text{ if }s\in(\overline{S},1].\end{cases}$ (9)
To understand Proposition 1, consider the auxiliary problem
$\max_{\widetilde{P}\in\Omega_{w}(\widetilde{P}^{*})}~{}~{}\int_{0}^{1}\overline{g}(s)\widetilde{P}(s)\,\mathrm{d}s,$ (10)
where the quantile virtual surplus $\widetilde{g}(s)$ in the original problem (7) is replaced by the ironed quantile virtual surplus $\overline{g}(s)$. For $s>\overline{S}$, the ironed virtual surplus is strictly negative, and thus the buyer should not trade. Consequently, any solution $\breve{P}$ of problem (10) must satisfy $\breve{P}(s)=0$. For $s\leq\overline{S}$; however, to maximize the objective, we would like to set $\widetilde{P}$ higher at points where $\overline{g}$ is larger. Since $\overline{g}$ is decreasing, this means that we want to set $\widetilde{P}(s)$ higher when $s$ is smaller. We see from (3) that no $\widetilde{P}\in\Omega_{w}(\widetilde{P}^{*})$ attains a larger value than $\widetilde{P}^{*}=(1-s)^{n-1}$ itself for small $s$; thus, $\widetilde{P}^{*}$ solves the auxiliary problem (10). Moreover, on each $\left[\underline{s}_{i},\overline{s}_{i}\right)$, since $\overline{G}$ is affine, $\overline{g}$ is constant, only the mean of the interim allocation on that interval matters. Therefore, letting
$\hat{P}(s)=\mathbb{E}\left[\widetilde{P}^{*}(t)~{}\big{|}~{}t\in\left[\underline{s}_{i},\overline{s}_{i}\right)\right]=\frac{\int_{\underline{s}_{i}}^{\overline{s}_{i}}(1-s)^{n-1}\,\mathrm{d}s}{\overline{s}_{i}-\underline{s}_{i}}$
on $\left[\underline{s}_{i},\overline{s}_{i}\right)$ for each $i\in\mathcal{I}$ and letting $\hat{P}=\widetilde{P}^{*}$ otherwise would make $\hat{P}$ a solution of (10). In fact, if $\hat{P}$ solves the auxiliary problem (10), it also solves the original problem if and only if it is constant on $\left[\underline{s}_{i},\overline{s}_{i}\right)$ for each $i\in\mathcal{I}$ [^8] This is the pooling property of Myerson (1981) and Toikka (2011), which states that for all open intervals $I\subseteq[0,1]$, $H(s)<\overline{H}(s)$ for all $s\in I$ implies that $\hat{P}$ must be constant on $I$. Therefore, interim allocation $\hat{P}$ defined by (9) solves problem (7), and thus is optimal.
Observe that ironing is required whenever the buyer’s virtual surplus function has an increasing region; that is, there exists an interval $(x,y)\subseteq[0,1]$ such that $g(q)$ is increasing on $(x,y)$. While it might be standard to assume that $F/f$ is increasing (and hence $-F/f$ is decreasing), requiring that $g(q)$ is decreasing on $[0,1]$ also imposes strong restrictions on $v(q)$: roughly speaking, the buyer’s marginal valuation of quality cannot be too high anywhere on the unit interval, [^9] If we further assume that both $v$ and $f$ are differentiable, $g(q)$ is decreasing if and only if $v^{\prime}(q)\leq 1+(F(q)/f(q))^{\prime}$ for all $q\in[0,1]$. For example, when qualities are uniformly distributed on $[0,1]$, $g(q)$ is decreasing if and only if $v^{\prime}(q)\leq 2$ for all $q\in[0,1]$. which implies that the buyer’s quality concerns cannot be too strong even locally. Therefore, in our problem, ironing might be necessary, due to the nature of the procurement setting.
@@ 3.2 Implementation: Bid-restricted auctions
To find the buyer-optimal procurement mechanism, we need to find some trading mechanisms that implement the optimal interim allocation rule $\hat{P}(s)$ identified in Proposition 1. To this end, we introduce the following class of mechanisms.
@@
A bid-restricted auction (BRA) is a sealed bid auction with $M\in\mathbb{N}$ bid intervals $\{[\underline{b}_{i},\overline{b}_{i}]\}_{i=1}^{M}$, where $\underline{b}_{1}\geq 0$, $\overline{b}_{M}\leq 1$, and for all $i=1,\ldots,M$, $\underline{b}_{i}\leq\overline{b}_{i}$ and $\overline{b}_{i}<\underline{b}_{i+1}$, with the following rules:
@@
  • The seller whose bid is the lowest wins the auction; in the event of a tie, the winning seller is chosen uniformly at random.
  • If the winning bid is the only bid in its bid interval $[\underline{b}_{i},\overline{b}_{i}]$ for some $i=1,\ldots,M$, and if the second-lowest bid equals $\underline{b}_{j}$ for some $j>i$ with $k$ other sellers bidding $\underline{b}_{j}$, [^10] If only one seller submits a bid, we take the second-lowest bid to be $\overline{b}_{M}$. then the winning seller receives a payment of $(\underline{b}_{j}+k\overline{b}_{j-1})/(k+1)$.
@@
Intuitively, a BRA is similar to a second-price auction, with two key exceptions: (i) sellers are allowed to bid only within certain bid intervals; and (ii) for any bid interval that is not the highest, there is a “payment reduction rule.” Under this rule, if the winning bid is the only bid in such a bid interval, the winning seller receives a payment that is determined by the other sellers’ bids, which may be lower than the second-lowest bid. The following result establishes an important property of BRAs; to simplify the statement, we adopt the convention that $\overline{b}_{0}:=0$.
@@
The structure inherited from a second-price auction and the “payment reduction rule” together imply that bidding as described in Lemma 3 is a weakly dominant strategy. Like a standard second-price auction, BRA also has the property that a seller’s bid affects the payment she receives only indirectly, through its influence on the identity of the winner. Moreover, when the payment received by the winner differs from the second-lowest bid, it helps eliminate the sellers’ incentives to take advantage of the “gaps” between the bid intervals.
To see this, consider a seller with quality $q\in[\overline{b}_{i-1},\underline{b}_{i})$. Because bids in the “gaps” are not allowed, it is more likely for someone to bid the lower bound of a bid interval. Therefore, the chance that some other sellers are also bidding $\underline{b}_{i}$ may not be negligible. In this case, if she bids $\underline{b}_{i}$ and it turns out to be the lowest bid, her chance of winning is at most 1/2, since ties are broken equiprobably. This may incentivize her to underbid—for example, bidding just below $\overline{b}_{i-1}$—to increase her chances of winning. This is where the payment reduction rule becomes relevant: the reduced payment must be chosen carefully to eliminate the incentive for underbidding. [^11] The reduced payment cannot be too low either, as doing so may encourage some sellers to overbid.
A function $h$ is said to be structured if there exists a finite partition of $[0,1]$ into intervals on which $h$ is either increasing or decreasing. If the virtual surplus $g$ is structured, then there exist $L\in\mathbb{N}$ and a collection of disjoint intervals $\{[\underline{s}_{i},\overline{s}_{i})\}_{i=1}^{L}$ such that $\overline{G}$ is affine on each interval $[\underline{s}_{i},\overline{s}_{i})$ for $i=1,\ldots,L$, and coincides with $G$ on the complement $[0,\overline{S}]\big{/}\bigcup_{i=1}^{L}\left[\underline{s}_{i},\overline{s}_{i}\right)$. We refer to the intervals $\{[\underline{s}_{i},\overline{s}_{i})\}_{i=1}^{L}$ as pooling intervals, and to the intervals comprising $[0,\overline{S}]\big{/}\bigcup_{i=1}^{L}\left[\underline{s}_{i},\overline{s}_{i}\right)$ as non-pooling intervals; it is easy to see that there are no more than $L+1$ non-pooling intervals.
@@
If the virtual surplus $g$ is structured, and $\overline{S}>0$, a buyer-optimal procurement mechanism is a BRA whose bid intervals are determined by
  • if $\underline{s}_{1}=0$, then $M=L$, $\underline{b}_{i}=F^{-1}(\overline{s}_{i})$ for all $i=1,\ldots,M$, $\overline{b}_{i}=F^{-1}(\underline{s}_{i+1})$ for $i=1,\ldots,M-1$, and $\overline{b}_{M}=F^{-1}(\overline{S})$;
  • if $\underline{s}_{1}>0$, then $M=L+1$, $\underline{b}_{1}=0$, $\underline{b}_{i}=F^{-1}(\overline{s}_{i-1})$ for $i=2,\ldots,M$, $\overline{b}_{i}=F^{-1}(\underline{s}_{i})$ for $i=1,\ldots,M-1$, and $\overline{b}_{M}=F^{-1}(\overline{S})$.
The optimality of BRA illustrates that prohibiting certain bids in an auction can alleviate quality concerns by eliminating vicious competition in prices, thereby “restoring” the auction’s optimality. The two conditions in Theorem 1 guarantee that a bid-restricted auction maps every non-pooling interval to a bid interval, and every pooling interval to a gap between two bid intervals. Consequently, every seller with quality quantile in a non-pooling interval bids her quality, and every seller with quality quantile in a pooling interval bids the highest quality of that interval.
By Lemma 3, every seller with quality quantile $s$ in a non-pooling interval wins if and only if the quality quantile of every other seller is above $s$, and therefore the interim allocation probability $\hat{P}(s)=(1-s)^{n-1}$. Furthermore, sellers with quality quantiles in the same pooling interval submit the same bid regardless of the exact quality quantiles, which indicates that the interim allocation probability is constant on the pooling intervals. Finally, if a seller has quality quantile above $\overline{S}$, she does not bid and hence never wins. In the proof, we formally show that the interim allocation induced by the BRA described in Theorem 1 is exactly the optimal interim allocation (9) in Proposition 1, which establishes the optimality of the described BRA.
@@
    (Manelli and Vincent, 1995)
    • If the virtual surplus is decreasing, the BRA with one bid interval $[0,F^{-1}(\overline{S})]$ is optimal, which is equivalent to a standard second-price auction with reserve price $F^{-1}(\overline{S})$.
    • If the virtual surplus is increasing, the BRA with a degenerate bid interval $\{1\}$ is optimal, which is equivalent to random allocation.
@@
    (Lopomo et al., 2023) If the virtual surplus is single-peaked, the BRA with one bid interval $[\underline{b},F^{-1}(\overline{S})]$ is optimal.

### v2

@@ 2.1 Transforming the problem: A reduced-form approach
Say that the collection of interim allocation probabilities $\{P_{s}\}_{s=1}^{n}$, where $P_{s}:[0,1]\to[0,1]$ for each $s$, is implementable if there exists a collection of ex-post allocation probabilities $\{p_{s}\}_{s=1}^{n}$ satisfying (F) that induces $\{P_{s}\}_{s=1}^{n}$ as its interim allocations; that is, for each $s=1,\ldots,n$ and all $q_{s}\in[0,1]$,
@@
Since the sellers are symmetric, it is without loss to restrict attention to symmetric interim allocations; thus, we can drop the subscript $s$ from $P_{s}$ and $q_{s}$ and write $P$ and $q$ instead. Consequently, the objective further reduces to
@@
As will become evident, introducing the quantile $s=F(q)$ makes identifying an optimal interim allocation straightforward. Define $\widetilde{P}(s)\coloneqq P(F^{-1}(s))$ as the quantile interim allocation; because $f>0$ on $(0,1]$, by Lemma 1, a collection of ex-post allocation probabilities is a part of an incentive-compatible mechanism if and only if the induced quantile interim allocation is decreasing.
Let $\mathcal{P}$ denote the set of decreasing functions mapping $[0,1]$ into itself. For any $\widetilde{P}^{\prime},\widetilde{P}^{\prime\prime}\in\mathcal{P}$, we say that $\widetilde{P}^{\prime}$ majorizes $\widetilde{P}^{\prime\prime}$, denoted by $\widetilde{P}^{\prime\prime}\prec\widetilde{P}^{\prime}$, if
$\int_{0}^{x}\widetilde{P}^{\prime\prime}(s)\,\mathrm{d}s\leq\int_{0}^{x}\widetilde{P}^{\prime}(s)\,\mathrm{d}s~~\text{ for all }x\in[0,1],$ (4)
with equality for $x=1$. Say that $\widetilde{P}^{\prime}$ weakly majorizes $\widetilde{P}^{\prime\prime}$, denoted by $\widetilde{P}^{\prime\prime}\prec_{w}\widetilde{P}^{\prime}$, if (4) holds but equality at $x=1$ is not required. Intuitively, $\widetilde{P}^{\prime}$ majorizes $\widetilde{P}^{\prime\prime}$ means that $\widetilde{P}^{\prime}$ is more variable, or more “spread out,” than $\widetilde{P}^{\prime\prime}$. In our context, this variability corresponds to the degree of discrimination among seller qualities: a trading mechanism that more sharply distinguishes between qualities generates a more variable interim allocation than one that pools qualities. While majorization keeps the ex-ante allocation probability $\int_{0}^{1}\widetilde{P}(s)\,\mathrm{d}s$ intact, weak majorization allows the ex-ante allocation probability to be strictly lower.
[…]
@@
A decreasing interim allocation probability $P$ is implementable if and only if the associated quantile interim allocation $\widetilde{P}(s)$ is weakly majorized by $\widetilde{P}^{*}$.
Border’s condition can be interpreted as follows: discriminating less by pooling qualities is always feasible, but discriminating more than the rule that always trades with the seller with the lowest quality is infeasible. Moreover, weak majorization allows for a reduction in the probability of trade through exclusion. [^9] Since $\int_{0}^{1}\widetilde{P}^{*}(s)\,\mathrm{d}s=1/n$, equality at $x=1$ in $\widetilde{P}\prec\widetilde{P}^{*}$ implies $n\int_{0}^{1}\widetilde{P}(s)\,\mathrm{d}s=1$, i.e., the buyer procures with probability one, whereas weak majorization allows this probability to be strictly below one.
@@
$h_{\gamma}(q)\coloneqq v(q)-q-\gamma\frac{F(q)}{f(q)}$
@@
$g(q)\coloneqq h_{1}(q)=v(q)-q-\frac{F(q)}{f(q)}$
[…]
@@
$\Omega_{w}(\widetilde{P}^{*})\coloneqq\left\{\widetilde{P}\in\mathcal{P}:\widetilde{P}\prec_{w}\widetilde{P}^{*}\right\}$.
In particular, the buyer-optimal interim allocation should solve the following problem:
@@
3 Buyer-optimal procurement mechanisms
For cleaner intuition and simpler notation, in this section we study a special case of our problem: the buyer’s expected payoff maximization problem, in which the welfare weight on the buyer is $\gamma=1$. In Section 3.1 we solve problem (6), and in Section 3.2 we identify a trading mechanism that implements the solution we find. All proofs in this section are relegated to Appendix B.
3.1 Buyer-optimal interim allocation
To ensure that the monotonicity constraint in problem (6) holds, the ironing technique (myerson1981, myerson1981; toikka2011, toikka2011) may be required. Define $\widetilde{g}(s)\coloneqq g(F^{-1}(s))$ as the (buyer’s) quantile virtual surplus, and let
$G(s)\coloneqq\int_{0}^{s}\widetilde{g}(x)\,\mathrm{d}x.$
Let $\overline{G}$ be the concave hull of $G$; that is, $\overline{G}(x)\coloneqq\sup\{y:(x,y)\in\mathrm{co}(G)\}$, where $\mathrm{co}(G)$ is the convex hull of the hypograph of $G$. Equivalently, $\overline{G}$ is the pointwise smallest concave function that lies above $G$. Call $\overline{g}\coloneqq\overline{G}^{\prime}$ the ironed quantile virtual surplus; [^11] Because $\overline{G}$ is concave, it is differentiable almost everywhere. At points where it is not differentiable, we define $\overline{g}$ as the right derivative. since $\overline{G}$ is concave, $\overline{g}$ is decreasing.
Proposition 1 characterizes a buyer-optimal interim allocation. Adopting the convention that $\sup\varnothing=0$, define
$\overline{S}\coloneqq\sup\{s\in[0,1]:\overline{g}(s)\geq 0\};$ (7)
in words, $\overline{S}$ is the highest quantile at which the ironed quantile virtual surplus is non-negative. Intuitively, $\overline{S}$ is the optimal exclusion cutoff: procuring from sellers with $s>\overline{S}$ is too expensive relative to the value they provide. Our assumptions on $v$ and $F$ guarantee that $\overline{S}$ is well-defined.
@@ 3.1 Buyer-optimal interim allocation
Let $\{\left[\underline{s}_{i},\overline{s}_{i}\right)\}_{i\in\mathcal{I}}$ denote a countable collection of disjoint intervals with $\left[\underline{s}_{i},\overline{s}_{i}\right)\subseteq[0,\overline{S}]$ for each $i\in\mathcal{I}$, such that
  • $\overline{G}$ is affine on $\left[\underline{s}_{i},\overline{s}_{i}\right)$ for each $i\in\mathcal{I}$; and
[…]
@@
$\hat{P}(s)=\begin{cases}(1-s)^{n-1}&\text{ if }s\in[0,\overline{S}]\big\backslash\bigcup_{i\in\mathcal{I}}\left[\underline{s}_{i},\overline{s}_{i}\right),\\ \frac{\int_{\underline{s}_{i}}^{\overline{s}_{i}}(1-s)^{n-1}\,\mathrm{d}s}{\overline{s}_{i}-\underline{s}_{i}}&\text{ if }s\in\left[\underline{s}_{i},\overline{s}_{i}\right),\\ 0&\text{ if }s\in(\overline{S},1].\end{cases}$ (8)
Proposition 1 follows from the results concerning maximizing linear functionals under a majorization constraint in kms. To better understand the intuition, consider first the simplest case in which the buyer’s virtual surplus $g$ is decreasing (and therefore so is the quantile virtual surplus $\widetilde{g}$). This implies that the buyer’s quality concerns are globally mild: roughly speaking, the buyer’s marginal valuation of quality is dominated by the marginal cost of quality everywhere on the unit interval. In this case, $G$ is concave, implying $G=\overline{G}$ and $\widetilde{g}=\overline{g}$; thus, ironing is not needed. For $s>\overline{S}$, $\widetilde{g}(s)$ is strictly negative, and thus the buyer would like to choose the minimal interim allocation 0 on $(\overline{S},1]$. For $s\leq\overline{S}$, the buyer would like to have a higher (lower) interim allocation when $\overline{g}$ is larger (smaller). Since $\overline{g}$ is decreasing, she would like to set $\widetilde{P}(s)$ as high (low) as possible when $s$ is small (large). This favors the most variable interim allocation $\widetilde{P}^{*}=(1-s)^{n-1}$, which corresponds to the buyer always procuring from the seller with the lowest quality.
When the buyer’s virtual surplus is not decreasing, $\widetilde{g}$ must be increasing around some quantile. This implies that quality concerns are locally severe: the value of procuring higher quality rises faster than the cost. Ideally, the buyer would prefer to favor higher-quality sellers in this region, but incentive compatibility requires the interim allocation probability to decrease with quality. Facing this conflict, the best the buyer can do is to make the interim allocation constant on an interval around that quantile. To satisfy feasibility (i.e., $\widetilde{P}\prec_{w}\widetilde{P}^{*}$), it suffices to set the interim allocation to the conditional mean of $\widetilde{P}^{*}$ on that interval. The ironing procedure identifies these intervals, which are the maximal intervals on which $G$ differs from $\overline{G}$; outside of these intervals, the logic in the previous paragraph applies, establishing the optimality of $\hat{P}$ as defined in (8). This is illustrated in Figure 2.
[…]
Figure 2: Illustration of Proposition 1 for $v(q)=-2q^{2}+4q$, $q\sim U[0,1]$ and $n=2$. In panel (a), the blue curve is $G$ and the orange curve is its concave hull, $\overline{G}$. In panel (b), the blue curve is $P^{*}(q)=1-q$ and the red curve is the optimal interim allocation, $\hat{P}(q)$.
@@ 3.2 Implementation: Bid-restricted auctions
To find the buyer-optimal procurement mechanism, we first need to find some trading mechanisms that implement the optimal interim allocation rule $\hat{P}(s)$ identified in Proposition 1. To this end, we introduce the following class of mechanisms.
@@
A bid-restricted auction (BRA) is a sealed-bid auction with $M\in\mathbb{N}$ bid intervals $\{[\underline{b}_{i},\overline{b}_{i}]\}_{i=1}^{M}$, where $\underline{b}_{1}\geq 0$, $\overline{b}_{M}\leq 1$, and for all $i=1,\ldots,M$, $\underline{b}_{i}\leq\overline{b}_{i}$ and $\overline{b}_{i}<\underline{b}_{i+1}$, with the following rules:
@@
  • The seller whose bid is the lowest wins the auction; ties are broken with equal probability among the sellers submitting the lowest bids.
  • If the winning bid is the only bid in its bid interval $\underline{b}_{i},\overline{b}_{i}]$ for some $i=1,\ldots,M$, and if the second-lowest bid equals $\underline{b}_{j}$ for some $j>i$ with a total of $k$ sellers who bid $\underline{b}_{j}$, [^12] If only one seller submits a bid, we take the second-lowest bid to be $\overline{b}_{M}$. then the winning seller receives a payment of $(\underline{b}_{j}+k\overline{b}_{j-1})/(k+1)$.
@@
Intuitively, a BRA is similar to a second-price auction, with two key exceptions: in a BRA, (i) sellers are allowed to bid only within certain bid intervals; and (ii) a payment reduction rule applies in specific cases. Under this rule, if the winning bid is the only bid in such a bid interval, the winning seller receives a payment that is determined by the other sellers’ bids, which may be lower than the second-lowest bid. The following result establishes an important property of BRAs; to simplify the statement, we adopt the convention that $\overline{b}_{0}:=0$.
@@
The structure inherited from a second-price auction and the payment reduction rule together imply that bidding as described in Lemma 3 is a weakly dominant strategy. As in standard second-price auctions, in BRAs a seller’s bid affects the payment she receives only indirectly, through its influence on the identity of the winner. In particular, except in a narrow set of contingencies, the winner is paid the second-lowest bid. Consequently, the standard second-price auction logic implies that if a seller’s quality lies in a bid interval $\underline{b}_{i},\overline{b}_{i})$, bidding her quality is weakly dominant. Likewise, if her quality lies in a gap $\overline{b}_{i-1},\,\underline{b}_{i})$, bidding above $\underline{b}_{i}$ can only reduce the probability of winning and cannot be profitable.
The only departure from the standard second-price payment is driven by the “gaps” between the bid intervals. To see this, consider a seller with quality $q\in[\overline{b}_{i-1},\,\underline{b}_{i})$. Under the strategy described in Lemma 3, the seller bids $\underline{b}_{i}$, and the same is true for any seller whose quality lies in $\overline{b}_{i-1},\,\underline{b}_{i})$; thus, ties at $\underline{b}_{i}$ may occur with positive probability. This gives the seller an incentive to “undercut across the gap” to increase her chances of winning. Consider a realized profile of opponents’ bids in which the lowest bid among the opponents equals $\underline{b}_{i}$ and exactly $k\geq 1$ opponents bid $\underline{b}_{i}$. If she bids $\underline{b}_{i}$, she wins with probability $1/(k+1)$, and when she wins her payment is $\underline{b}_{i}$, so her expected payoff in this contingency is $(\underline{b}_{i}-q)/(k+1)$. Submitting a bid $b\leq\overline{b}_{i-1}$ instead makes her the unique lowest bidder in the contingency above. Under a standard second-price payment, she would still be paid $\underline{b}_{i}$, which would yield payoff $\underline{b}_{i}-q>(\underline{b}_{i}-q)/(k+1)$. The payment reduction rule lowers the winner’s payment exactly in these contingencies, eliminating this underbidding incentive and restoring the weak dominance of the strategy described in Lemma 3.
To pin down the reduced payment $p$ in such cases, we need to ensure that
$p-q\leq\frac{\underline{b}_{i}-q}{k+1}\quad\text{for all }q\in[\overline{b}_{i-1},\underline{b}_{i}),$
or equivalently,
$p\leq\frac{\underline{b}_{i}+kq}{k+1}\quad\text{for all }q\in[\overline{b}_{i-1},\underline{b}_{i}).$
Because the right-hand side is strictly increasing in $q$, it suffices that the inequality holds at $q=\overline{b}_{i-1}$, i.e., $p\leq(\underline{b}_{i}+k\overline{b}_{i-1})/(k+1)$. Conversely, the reduced payment cannot be chosen strictly below this value: if $p<(\underline{b}_{i}+k\overline{b}_{i-1})/(k+1)$, there exists $\varepsilon>0$ such that a seller with quality $q\in(\overline{b}_{i-1}-\varepsilon,\overline{b}_{i-1})$ strictly prefers to overbid to $\underline{b}_{i}$ in the contingency above. Therefore, weak dominance pins down the reduced payment uniquely as $p=(\underline{b}_{i}+k\overline{b}_{i-1})/(k+1)$.
[…]
@@
Suppose the buyer’s virtual surplus $g$ is structured. A buyer-optimal procurement mechanism is a BRA with bid intervals $\{[\underline{b}_{i},\overline{b}_{i}]\}_{i=1}^{M}$ defined as follows:
  • If $\underline{s}_{1}=0$, then $M=L$ and
    $\underline{b}_{i},\overline{b}_{i}]=\begin{cases}\big[F^{-1}(\overline{s}_{i}),\,F^{-1}(\underline{s}_{i+1})\big],&i=1,\ldots,L-1,\\[2.0pt] \big[F^{-1}(\overline{s}_{L}),\,F^{-1}(\overline{S})\big],&i=L.\end{cases}$
  • If $\underline{s}_{1}>0$, then $M=L+1$ and
    $\underline{b}_{i},\overline{b}_{i}]=\begin{cases}\big[0,\,F^{-1}(\underline{s}_{1})\big],&i=1,\\[2.0pt] \big[F^{-1}(\overline{s}_{i-1}),\,F^{-1}(\underline{s}_{i})\big],&i=2,\ldots,L,\\[2.0pt] \big[F^{-1}(\overline{s}_{L}),\,F^{-1}(\overline{S})\big],&i=L+1.\end{cases}$
Although a standard second-price procurement auction is generally not optimal in our setting due to quality concerns, Theorem 1 shows that restricting bids to suitably chosen intervals renders a simple second-price-like auction format optimal. The construction in Theorem 1 is chosen so that each non-pooling interval of the optimal interim allocation corresponds to a bid interval, while each pooling interval corresponds to a gap between two bid intervals. Consequently, a seller with a quality quantile in a non-pooling interval bids her quality as in a standard second-price auction, and a seller with a quality quantile in a pooling interval bids the highest quality of that interval.
Lemma 3 implies that if a seller’s quality quantile $s$ is in a non-pooling interval, she wins if and only if her quality quantile is the lowest; i.e., every other seller’s quality quantile exceeds $s$. Therefore, her quantile interim allocation equals $\hat{P}(s)=(1-s)^{n-1}$ on non-pooling intervals. Furthermore, sellers with quality quantiles in the same pooling interval submit the same bid regardless of their exact quality quantiles, which indicates that the interim allocation probability is constant on that interval. Finally, if a seller has a quality quantile above $\overline{S}$, she does not bid and hence never wins. In the proof, we formally show that the interim allocation induced by the BRA described in Theorem 1 coincides with the optimal interim allocation (8) in Proposition 1, which establishes its optimality.
@@
    • If the virtual surplus is decreasing, the BRA with one bid interval $0,F^{-1}(\overline{S})]$ is optimal, which is equivalent to a standard second-price auction with reserve price $F^{-1}(\overline{S})$.
    • If the virtual surplus is increasing, the BRA with a degenerate bid interval $\{1\}$ is optimal, which is equivalent to selecting the winning seller at random.
@@
    (lpv) If the virtual surplus is single-peaked, the BRA with one bid interval $\underline{b},F^{-1}(\overline{S})]$ is optimal.